I have been asked to find the domain of these two functions:
f(x)=x^2/x^2-2 and g(x)=x+4/x^3-4x
a) the domain of f(x) in interval form: x cannot be -1 or 1
b)the domain of g(x) in interval form: x cannot be -2, 0, or 2
How do I write these in interval notation? This is my best guess, would this be correct?
a)(-infinity, -1)U(1,infinity)
b)(-infinity,-2)U(0)U(2,infinity)

Answer by stanbon(75887)

find the domain of these two functions:
f(x)=x^2/x^2-2 and g(x)=x+4/x^3-4x
a) the domain of f(x) in interval form: x cannot be -sqrt(2) or +sqrt(2)
(-inf,-sqrt(2))U(-sqrt(2),+sqrt(2))U(sqrt(2),+inf)
-----------------------------
b)the domain of g(x) in interval form: x cannot be -2, 0, or 2
How do I write these in interval notation?
(-infinity,-2)U(-2,0)U(0,2)U(2,infinity)
=======================
Cheers,
Stan H.
===============